OpenStudy (anonymous):

If x^2 + y^2 = 25 and dy/dt = 4, find dx/dt when y = 3

OpenStudy (freckles):

So @jsa are you having trouble differentiating w.r.t t?

OpenStudy (anonymous):

yes :/ help?

OpenStudy (freckles):

\[\frac{d}{dt}x^2=2x \frac{dx}{dt} \text{ by chain rule } \]

OpenStudy (freckles):

you try differentiating y^2 w.r.t t

OpenStudy (freckles):

@jsa have you tried yet?

OpenStudy (anonymous):

I'm really confused...

OpenStudy (freckles):

On chain rule?

OpenStudy (freckles):

do you have any problems with this: \[\frac{d}{dx}x^2=2x ?\]